OpenStudy (anonymous):

Price of a ticket to the red sox games rises from $10 to $12 per game. The quantity of tickets sold falls 160000 to 144000. Calculate the elasticity of demand, whether elastic, inelastic or unit elastic. Will revenue rise, decline or stay the same? Calculus

OpenStudy (anonymous):

elasticity: (%change in demand)/(%change in price) when price changes from 10 to 12, demand falls from 160000 to 144000 (%change in price) = (12-10)/10 = 0.2 (%change in demand) = (144000 - 160000)/160000 = -0.1 elasticity = -0.1/0.2 = -0.5 the elasticity is < than 1 so Demand is price inelastic

OpenStudy (anonymous):

do you know how to find the revenue of this problem?

OpenStudy (anonymous):

yes revenue = price times product sold old revenue = 10 times 160000 new revenue = 12 times 144000 then calculate the difference

OpenStudy (anonymous):

ok thanks